### v1

# SwiftEmbed: Ultra-Fast Text Embeddings via Static Token Lookup for Real-Time Applications

Abstract
We present a static token lookup methodology for text embedding generation that achieves 1.12 ms p50 latency for single text embeddings while maintaining 60.6 MTEB average score across 8 representative tasks, corresponding to 89% of contextual model quality. The Rust implementation delivers 50,000 requests per second throughput through static embedding lookup, optimized mean pooling, and zero-copy IEEE754 binary serialization. Evaluation demonstrates exceptional duplicate detection performance (90.1% AP), strong semantic similarity (76.1% Spearman correlation), and domain-specific performance ranging from 75% to 131% of baseline across specialized domains. The system enables real-time embedding applications where sub-5ms latency is critical.
1 Introduction
Text embeddings have become fundamental to natural language processing applications, yet transformer-based approaches introduce prohibitive latency for real-time scenarios requiring sub-millisecond response times. While models like BERT [3] achieve state-of-the-art semantic quality, their multi-layer attention mechanisms create computational bottlenecks that limit deployment in latency-sensitive environments.
This work introduces a paradigm shift by replacing transformer inference with static token lookup, leveraging pre-computed token-level embeddings from compact models combined with SIMD-optimized aggregation. Our approach achieves 1.12ms response times and 50,000 RPS throughput while maintaining 89% of contextual model quality, enabling previously impossible real-time applications.
The contributions include: a static token lookup methodology eliminating transformer inference; SIMD and zero-copy optimizations achieving 30-50% memory reduction; a production-ready Rust implementation supporting 10,000+ concurrent connections; and comprehensive empirical analysis of speed-quality trade-offs.
2 Related Work
2.1 Transformer-Based Embedding Models
Transformer architectures from BERT [3] to Sentence-BERT [9] have established contemporary standards for semantic representations through multi-layer attention mechanisms [13]. While achieving superior quality, their computational complexity renders them unsuitable for ultra-low latency scenarios. Compression techniques like DistilBERT [11] and TinyBERT [5] reduce model sizes but retain full transformer inference, fundamentally limiting latency improvements. Recent advances in efficient transformers [21] and model compression [20] have explored various optimization paths while maintaining transformer architecture.
2.2 Efficient Embedding Systems
Systems like Sentence Transformers [10], Txtai [8], and Weaviate [14] optimize around transformer inference rather than eliminating it. FastText [1] provides subword embeddings but lacks contextual understanding, while traditional approaches like GloVe [17] and Word2Vec [18] offer static embeddings without modern contextual awareness. Our methodology occupies a novel position by completely bypassing transformer inference while maintaining competitive semantic performance through optimized aggregation of pre-trained token representations. Recent work on quantization [12, 15] and efficient retrieval [35, 29] has addressed specific aspects of efficiency, but none achieve the comprehensive latency reductions demonstrated in our approach.
3 Methodology
3.1 Computational Complexity Analysis
The computational complexity analysis reveals our fundamental advantage. Traditional transformers with $L$ layers, hidden dimension $d_{h}$, and sequence length $n$ exhibit quadratic complexity $\mathcal{C}_{\text{transformer}}=O(L\cdot n^{2}\cdot d_{h}+L\cdot n\cdot d_{h}^{2})$, while our static approach reduces this to linear complexity $\mathcal{C}_{\text{static}}=O(n+d)$.
The theoretical speedup factor $\gamma=\frac{\mathcal{C}_{\text{transformer}}}{\mathcal{C}_{\text{static}}}\approx O(L\cdot n\cdot d_{h})$ explains the observed 20× empirical improvements. For typical parameters ($L=12$, $d_{h}=768$, $n=512$), $\gamma\approx 4.7\times 10^{6}$ accounts for the dramatic performance gains.
3.2 Mathematical Formulation
For vocabulary $\mathcal{V}$ and embedding matrix $E\in\mathbb{R}^{|\mathcal{V}|\times d}$, text sequence $S=(w_{1},w_{2},\ldots,w_{n})$ undergoes tokenization $\tau(S)=(t_{1},t_{2},\ldots,t_{m})$, embedding lookup $\mathbf{e}_{i}=E[\text{index}(t_{i})]$, attention-weighted mean pooling $\mathbf{h}=\frac{\sum_{i=1}^{m}a_{i}\cdot\mathbf{e}_{i}}{\sum_{i=1}^{m}a_{i}}$, and L2 normalization $\mathbf{f}=\frac{\mathbf{h}}{\|\mathbf{h}\|_{2}}\in\mathbb{S}^{d-1}$.
Batch processing extends this through concatenated token sequences and parallel aggregation, maintaining linear complexity while maximizing hardware utilization. The formulation draws inspiration from traditional static embeddings [17, 18] while incorporating modern optimization techniques.
3.3 System Architecture and Optimizations
The Rust implementation [6] with Axum framework [2] achieves 8% higher throughput through superior hyper/tokio integration. Key optimizations include static embedding lookup via single tensor operations, SIMD JSON processing with 256-bit vector instructions, memory prefetching reducing cache misses by 30-50%, and zero-copy binary serialization eliminating memory copies. The Candle framework [4] provides efficient tensor operations and automatic backend selection.
Three response formats cater to different needs: JSON for compatibility, binary IEEE754 for maximum performance, and JSONL for streaming scenarios. The architecture incorporates lessons from efficient inference systems [34] and high-performance computing principles [29].
4 Experimental Evaluation
4.1 Experimental Setup
Our evaluation employs wrk with specialized Lua scripts for single text processing, batch processing, JSON Lines streaming, and variable batch sizes. Configuration uses 12 threads with 400 concurrent connections over 30-second test durations. We evaluate on 8 representative MTEB tasks [19] covering classification, clustering, retrieval, similarity, and pair classification, comparing against Sentence-BERT, GTE-tiny, BGE-micro, TensorRT-optimized transformers, and traditional static embeddings. Additional evaluation includes BEIR [22] for zero-shot retrieval and established benchmarks [23, 24, 25] for comprehensive assessment.
4.2 Comprehensive Quality Evaluation
Table 1: MTEB Benchmark Results (8 representative tasks)
Method Classification Clustering Retrieval STS Pair Class.
(Acc) (V-measure) (nDCG@10) (Spearman) (AP)
Sentence-BERT 75.2 42.3 51.4 85.2 84.7
GTE-tiny 72.1 40.8 48.7 82.3 82.1
BGE-micro 71.4 39.6 47.2 81.8 81.4
Jina-v2-small 73.8 41.2 49.6 83.1 83.2
DistilBERT 70.3 38.4 45.8 79.6 79.8
FastText 58.2 28.4 32.1 65.3 68.2
GloVe-840B 56.7 27.1 30.8 62.4 65.3
Our Method 58.9 35.6 42.1 76.1 90.1
Our method achieves 60.6 MTEB average score, corresponding to 89% of Sentence-BERT quality while providing substantial performance advantages. The evaluation reveals exceptional performance on duplicate detection tasks with 90.1% average precision on SprintDuplicateQuestions, outperforming transformer baselines. Retrieval tasks maintain 82% of BERT performance (nDCG@10: 42.1 vs 51.4 on ArguAna), while semantic similarity demonstrates 89% correlation strength (0.761 vs 0.852 Spearman average). Classification shows domain-dependent characteristics with strong Banking77 performance (72.7%) but limitations on EmotionClassification (45.2%).
4.3 Performance Analysis
Table 2: Performance Comparison with State-of-the-Art Methods
Method Model Size p50 Lat. p99 Lat. Throughput Memory Quality
(MB) (ms) (ms) (RPS) (GB) (MTEB)
Sentence-BERT 440 45 120 2,500 1.8 67.8
GTE-tiny 133 32 85 3,800 0.8 65.2
BGE-micro 98 25 68 4,500 0.6 64.3
TensorRT-BERT 440 12 35 8,500 2.2 67.8
Quantized BERT 110 15 40 7,200 0.5 65.8
FastText 2.5 8 15 15,000 0.1 50.4
GloVe-840B 5.4 6 12 18,000 0.1 48.5
Our Method 32 1.12 5.04 50,000 0.2 60.6
Our approach demonstrates 20× improvement over Sentence-BERT in single request throughput and 8× improvement over TensorRT-BERT in single request latency. The 32MB model footprint enables deployment in resource-constrained environments, with runtime memory requirements of 0.2GB versus 1.8GB for Sentence-BERT. The combination achieves 3.3× improvement over FastText++ while maintaining 14% higher quality, with linear scaling characteristics versus quadratic degradation in transformer-based approaches.
4.4 Ablation Studies
Table 3: Static Embedding Model Comparison
Embedding Model Vocab Size Dim Model Size MTEB Avg Latency
(MB) Score (ms)
GloVe-50d 400k 50 77 42.3 1.2
GloVe-300d 840k 300 989 48.5 1.8
Word2Vec-300d 3M 300 3,500 45.7 2.1
FastText-300d 2M 300 2,400 50.4 1.9
FastText-subword 2M 300 2,400 52.1 2.3
Potion-base-8M 30k 384 32 56.3 1.7
Potion-base-4M 30k 256 24 54.8 1.5
Potion-base-8M achieves 16% higher MTEB score than GloVe-840B with 97% smaller model size, providing optimal quality-size trade-off. Subword models add 21% latency overhead with minimal quality gain, while smaller 30k vocabularies enable better cache utilization compared to 840k vocabularies.
Table 4: Impact of Embedding Dimension
Dimension MTEB Avg Latency Memory Cache Hits Bandwidth
Score (ms) (MB) (%) (GB/s)
128 52.1 0.8 12 98.2 2.1
256 54.8 1.2 24 96.5 3.8
384 56.3 1.7 32 94.1 5.2
512 56.8 2.3 48 89.3 6.9
768 57.2 3.5 72 82.1 9.8
The 384-dimensional configuration provides optimal quality-performance trade-off, balancing 56.3 MTEB score with 1.7ms latency and 32MB memory footprint while maintaining 94.1% cache hit rate.
4.5 Throughput Scalability Analysis
Table 5: Throughput Scalability Comparison
Method Single Batch-10 Batch-100 Max Conc. Scaling
RPS RPS RPS
Sentence-BERT 2,500 1,200 250 500 Quadratic
DistilBERT 4,200 2,100 420 800 Quadratic
TensorRT-BERT 8,500 4,000 850 1,000 Sub-quadratic
FastText++ 15,000 12,000 8,000 2,000 Linear
Our Method 50,000 10,000 2,000 10,000+ Linear
Our approach demonstrates superior scalability with linear characteristics versus quadratic degradation in transformer-based methods. The system maintains 50,000 RPS for single requests and scales to support 10,000+ concurrent connections, enabling high-density deployment scenarios.
4.6 Domain-Specific Performance
Table 6: Domain-Specific Evaluation Results
Domain Similarity Semantic Clustering Avg Latency Relative
Correlation Coherence ARI (ms) Performance
General 0.564 0.210 0.207 0.45 Baseline
Medical 0.422 0.221 0.175 0.47 75%
Legal 0.533 0.343 0.135 0.48 95%
Technical 0.563 0.278 0.059 0.51 100%
Social Media 0.606 0.276 0.029 0.44 108%
Scientific 0.739 0.326 -0.027 0.52 131%
Scientific text demonstrates strongest performance with 131% relative effectiveness, attributed to consistent terminology and limited polysemy. Legal text performs well (95%) due to formal language structures, while medical text shows reduced effectiveness (75%) from specialized vocabulary and contextual dependencies. Social media content achieves 108% performance, benefiting from standardized expressions and hashtag usage.
4.7 Downstream Task Performance
Table 7: Downstream Task Evaluation
Task Our Score Baseline Performance Speed-Quality
(MiniLM) Ratio Trade-off
Semantic Search 0.877 0.947 0.926 Acceptable
Classification 0.500 0.667 0.750 Acceptable
Clustering 0.627 0.792 0.792 Good
Duplicate Detection 1.000 1.000 1.000 Optimal
Downstream evaluation reveals 92.6% performance retention for semantic search and 100% for duplicate detection compared to compact transformer baselines. The approach demonstrates optimal trade-offs for duplicate detection applications while maintaining acceptable performance degradation for classification and clustering tasks within sub-2ms latency constraints.
4.8 Sequence Length Impact Analysis
Table 8: Performance vs Sequence Length Characteristics
Sequence Length Avg Latency Avg Text Similarity Embedding Relative
Category (ms) Length Score Norm Speed
Short (1-20 words) 0.453 20 1.000 1.000 Baseline
Medium (20-50) 0.468 102 1.000 1.000 97%
Long (50-100) 0.699 255 1.000 1.000 65%
Very Long (100+) 1.242 510 1.000 1.000 36%
The system demonstrates sub-linear latency scaling from 0.45ms for short texts to 1.24ms for very long texts, maintaining consistent similarity scores and embedding norms across sequence lengths. This scaling characteristic ensures predictable performance for diverse text inputs while preserving semantic quality.
4.9 Multilingual Performance
Table 9: Multilingual Performance Analysis
Language Cross-Language Within-Language Embedding Relative
Similarity Coherence Norm Performance
English 1.000 0.359 1.000 Baseline
Spanish 0.225 0.468 1.000 22.5%
French 0.198 0.271 1.000 19.8%
German 0.170 0.405 1.000 17.0%
Multilingual analysis shows significant degradation, achieving only 17-23% of English performance on other languages. This indicates primary optimization for English, with cross-language similarity scores of 0.225 for Spanish, 0.198 for French, and 0.170 for German compared to the English baseline.
4.10 Failure Analysis
Table 10: Qualitative Failure Analysis
Input Text BERT Similarity Our Similarity Failure Type
”The bank approved the loan” vs ”Financial institution granted credit” 0.92 0.78 Semantic understanding
”Apple released iOS” vs ”The fruit company launched software” 0.85 0.42 Entity disambiguation
”bank” (financial) vs ”bank” (river) 0.15 0.95 Polysemy handling
”He didn’t not go” vs ”He went” 0.88 0.31 Double negation
”The cat sat on the mat” vs ”On the mat sat the cat” 0.96 0.72 Word order sensitivity
Systematic failure analysis reveals consistent patterns: polysemy handling accounts for 35% of failures, compositional semantics 28%, named entities 22%, and negation/modality 15%. The approach struggles with contextual disambiguation, particularly for words with multiple meanings and phrases where meaning differs from individual word combinations.
5 Theoretical Analysis
5.1 Speed-Quality Trade-off Formulation
From an information-theoretic perspective, traditional transformers maximize mutual information $I_{\text{transformer}}(X;Y)=H(Y)-H(Y|X)\approx\log_{2}(d_{h})-\epsilon_{\text{contextual}}$, while our static approach achieves $I_{\text{static}}(X;Y)=H(Y)-H(Y|X)\approx\log_{2}(d)-\epsilon_{\text{static}}$ with $\epsilon_{\text{static}}>\epsilon_{\text{contextual}}$ due to contextual information loss.
The approximation error is bounded by $\|\mathbf{f}_{\text{static}}-\mathbf{f}_{\text{transformer}}\|_{2}\leq\sqrt{\frac{2}{d}}\times\text{Context\_variance}+\delta_{\text{pooling}}$, where $\delta_{\text{pooling}}$ represents error from mean pooling versus attention mechanisms.
5.2 Memory Efficiency Analysis
Our static approach exhibits linear memory scaling $\text{Memory}_{\text{static}}(B,n,d)=|\mathcal{V}|\times d\times 4+B\times n\times d\times 4+O(1)$ compared to transformer requirements $\text{Memory}_{\text{transformer}}(B,n,d_{h},L)=O(L\times d_{h}^{2}+B\times n^{2}\times d_{h})$.
The memory efficiency gain $\rho=\frac{\text{Memory}_{\text{transformer}}}{\text{Memory}_{\text{static}}}=\frac{L\times d_{h}^{2}+B\times n^{2}\times d_{h}}{|\mathcal{V}|\times d+B\times n\times d}$ yields $\rho\approx 0.73$ for typical parameters, explaining the observed memory reductions.
5.3 Serialization Performance
Table 11: Serialization Format Performance Characteristics
Format Complexity Serialization Memory Bandwidth
Time Overhead Efficiency
Binary $O(B\times d)$ 0.1ms 0% 100%
JSON (SIMD) $O(B\times d\times\log d)$ 0.5ms 15-20% 60-70%
JSONL $O(B\times d\times\log d)$ 0.3ms 8-12% 75-85%
Binary IEEE754 format achieves zero-copy performance through direct memory mapping, eliminating serialization overhead entirely. SIMD-optimized JSON processing provides compatibility with 2.5-3.2× speedup, while JSONL balances performance and streaming capabilities.
6 Discussion
6.1 Technical Design Justifications
Table 12: Language Comparison for Implementation
Language Zero-Copy Memory Safety SIMD Support Async Perf. Build Size
Support (M req/s) (MB)
Rust Native Compile-time Excellent 2.8 12
C++ Manual Runtime checks Excellent 2.6 18
Go Limited GC-based Limited 1.9 25
Python None GC-based Via libraries 0.3 150+
Rust provides native zero-copy support, compile-time memory safety, and excellent SIMD support, achieving 2.8 million requests per second in asynchronous performance. The language selection enables predictable latency without GC pauses and direct access to CPU vector instructions, drawing from modern systems programming principles [6].
6.2 Application Scope and Limitations
Our method excels in real-time search and retrieval systems requiring sub-5ms embedding generation, high-throughput processing pipelines, resource-constrained edge deployment, and duplicate detection applications where it outperforms contextual approaches. The 32MB model footprint and linear scaling characteristics make it suitable for high-density server configurations and limited hardware environments.
However, semantic degradation of 5-15% compared to contextual models represents a fundamental trade-off, with limitations in contextual disambiguation, multilingual performance achieving only 17-23% of English effectiveness, and domain specificity requiring careful application selection. These limitations align with known challenges in static embedding approaches [1, 17] while providing unprecedented latency benefits.
7 Conclusion
This research demonstrates that static token lookup achieves sub-2ms response times and 50,000 RPS throughput while maintaining 60.6 MTEB average score (89% of Sentence-BERT quality). The approach provides exceptional duplicate detection performance (90.1% AP), strong semantic similarity capabilities (76.1% Spearman correlation), and domain-specific effectiveness ranging from 75% to 131% of baseline.
The 20× efficiency improvement over traditional approaches enables new real-time applications while reducing computational requirements. Future work includes hybrid contextualization frameworks [33], adaptive quantization techniques [12, 15], and cross-modal static embeddings to further bridge computational efficiency and semantic precision.
Limitations of Evaluation
We acknowledge evaluation on 8 representative MTEB tasks rather than the full benchmark, performance comparisons based on published results, theoretical ablation studies, proprietary implementation, and primary optimization for English with multilingual limitations. The benchmarking infrastructure [16] enables verification of performance characteristics while the core implementation remains proprietary.
Broader Impact
Our approach enables more environmentally sustainable NLP applications through dramatically reduced computational requirements, contributing to reduced computational carbon footprint. The efficiency improvements align with growing concerns about AI environmental impact while maintaining practical utility for real-world applications.
Acknowledgments
We thank the developers of open-source tools including Rust [6], Axum [2], Candle [4], and the broader open-source community for enabling this research.

### v2

@@
-SwiftEmbed: Ultra-Fast Text Embeddings via Static Token Lookup for Real-Time Applications
+
 Abstract
@@
 The system enables real-time embedding applications where sub-5ms latency is critical.
+keywords:
+Ultra-low latency, text embeddings, Static token lookup, real-time applications, Zero-copy serialization, Edge deployment
 1 Introduction
 Text embeddings have become fundamental to natural language processing applications, yet transformer-based approaches introduce prohibitive latency for real-time scenarios requiring sub-millisecond response times.
-While models like BERT [3] achieve state-of-the-art semantic quality, their multi-layer attention mechanisms create computational bottlenecks that limit deployment in latency-sensitive environments.
+While models like BERT [devlin2018bert] achieve state-of-the-art semantic quality, their multi-layer attention mechanisms create computational bottlenecks that limit deployment in latency-sensitive environments.
 This work introduces a paradigm shift by replacing transformer inference with static token lookup, leveraging pre-computed token-level embeddings from compact models combined with SIMD-optimized aggregation.
@@
 2.1 Transformer-Based Embedding Models
-Transformer architectures from BERT [3] to Sentence-BERT [9] have established contemporary standards for semantic representations through multi-layer attention mechanisms [13].
+Transformer architectures from BERT [devlin2018bert] to Sentence-BERT [reimers2019sentence] have established contemporary standards for semantic representations through multi-layer attention mechanisms [vaswani2017attention].
 While achieving superior quality, their computational complexity renders them unsuitable for ultra-low latency scenarios.
-Compression techniques like DistilBERT [11] and TinyBERT [5] reduce model sizes but retain full transformer inference, fundamentally limiting latency improvements.
-Recent advances in efficient transformers [21] and model compression [20] have explored various optimization paths while maintaining transformer architecture.
+Compression techniques like DistilBERT [sanh2019distilbert] and TinyBERT [jiao2019tinybert] reduce model sizes but retain full transformer inference, fundamentally limiting latency improvements.
+Recent advances in efficient transformers [tay2023transformer] and model compression [xiao2023survey] have explored various optimization paths while maintaining transformer architecture.
 2.2 Efficient Embedding Systems
-Systems like Sentence Transformers [10], Txtai [8], and Weaviate [14] optimize around transformer inference rather than eliminating it.
-FastText [1] provides subword embeddings but lacks contextual understanding, while traditional approaches like GloVe [17] and Word2Vec [18] offer static embeddings without modern contextual awareness.
+Systems like Sentence Transformers [reimers2019sentence2], Txtai [neuml2020txtai], and Weaviate [weaviate2019] optimize around transformer inference rather than eliminating it.
+FastText [bojanowski2017enriching] provides subword embeddings but lacks contextual understanding, while traditional approaches like GloVe [pennington2014glove] and Word2Vec [mikolov2013efficient] offer static embeddings without modern contextual awareness.
 Our methodology occupies a novel position by completely bypassing transformer inference while maintaining competitive semantic performance through optimized aggregation of pre-trained token representations.
-Recent work on quantization [12, 15] and efficient retrieval [35, 29] has addressed specific aspects of efficiency, but none achieve the comprehensive latency reductions demonstrated in our approach.
+Recent work on quantization [shen2020qbert, zafrir2019q8bert] and efficient retrieval [khattab2020colbert, johnson2019billion] has addressed specific aspects of efficiency, but none achieve the comprehensive latency reductions demonstrated in our approach.
 3 Methodology
@@
 Batch processing extends this through concatenated token sequences and parallel aggregation, maintaining linear complexity while maximizing hardware utilization.
-The formulation draws inspiration from traditional static embeddings [17, 18] while incorporating modern optimization techniques.
+The formulation draws inspiration from traditional static embeddings [pennington2014glove, mikolov2013efficient] while incorporating modern optimization techniques.
 3.3 System Architecture and Optimizations
-The Rust implementation [6] with Axum framework [2] achieves 8% higher throughput through superior hyper/tokio integration.
+The Rust implementation [klabnik2018rust] with Axum framework [tokio2021axum] achieves 8% higher throughput through superior hyper/tokio integration.
 Key optimizations include static embedding lookup via single tensor operations, SIMD JSON processing with 256-bit vector instructions, memory prefetching reducing cache misses by 30-50%, and zero-copy binary serialization eliminating memory copies.
-The Candle framework [4] provides efficient tensor operations and automatic backend selection.
+The Candle framework [huggingface2023candle] provides efficient tensor operations and automatic backend selection.
 Three response formats cater to different needs: JSON for compatibility, binary IEEE754 for maximum performance, and JSONL for streaming scenarios.
-The architecture incorporates lessons from efficient inference systems [34] and high-performance computing principles [29].
+The architecture incorporates lessons from efficient inference systems [lee2019latency] and high-performance computing principles [johnson2019billion].
 4 Experimental Evaluation
@@
 Configuration uses 12 threads with 400 concurrent connections over 30-second test durations.
-We evaluate on 8 representative MTEB tasks [19] covering classification, clustering, retrieval, similarity, and pair classification, comparing against Sentence-BERT, GTE-tiny, BGE-micro, TensorRT-optimized transformers, and traditional static embeddings.
-Additional evaluation includes BEIR [22] for zero-shot retrieval and established benchmarks [23, 24, 25] for comprehensive assessment.
+We evaluate on 8 representative MTEB tasks [muennighoff2022mtEB] covering classification, clustering, retrieval, similarity, and pair classification, comparing against Sentence-BERT, GTE-tiny, BGE-micro, TensorRT-optimized transformers, and traditional static embeddings.
+Additional evaluation includes BEIR [thakur2021beir] for zero-shot retrieval and established benchmarks [wang2018glue, rajpurkar2016squad, conneau2018xnli] for comprehensive assessment.
 4.2 Comprehensive Quality Evaluation
@@
 Rust provides native zero-copy support, compile-time memory safety, and excellent SIMD support, achieving 2.8 million requests per second in asynchronous performance.
-The language selection enables predictable latency without GC pauses and direct access to CPU vector instructions, drawing from modern systems programming principles [6].
+The language selection enables predictable latency without GC pauses and direct access to CPU vector instructions, drawing from modern systems programming principles [klabnik2018rust].
 6.2 Application Scope and Limitations
@@
 However, semantic degradation of 5-15% compared to contextual models represents a fundamental trade-off, with limitations in contextual disambiguation, multilingual performance achieving only 17-23% of English effectiveness, and domain specificity requiring careful application selection.
-These limitations align with known challenges in static embedding approaches [1, 17] while providing unprecedented latency benefits.
+These limitations align with known challenges in static embedding approaches [bojanowski2017enriching, pennington2014glove] while providing unprecedented latency benefits.
 7 Conclusion
@@
 The 20× efficiency improvement over traditional approaches enables new real-time applications while reducing computational requirements.
-Future work includes hybrid contextualization frameworks [33], adaptive quantization techniques [12, 15], and cross-modal static embeddings to further bridge computational efficiency and semantic precision.
+Future work includes hybrid contextualization frameworks [lewis2020retrieval], adaptive quantization techniques [shen2020qbert, zafrir2019q8bert], and cross-modal static embeddings to further bridge computational efficiency and semantic precision.
 Limitations of Evaluation
 We acknowledge evaluation on 8 representative MTEB tasks rather than the full benchmark, performance comparisons based on published results, theoretical ablation studies, proprietary implementation, and primary optimization for English with multilingual limitations.
-The benchmarking infrastructure [16] enables verification of performance characteristics while the core implementation remains proprietary.
+The benchmarking infrastructure [lansiaux2025benchmarks] enables verification of performance characteristics while the core implementation remains proprietary.
 Broader Impact
@@
 Acknowledgments
-We thank the developers of open-source tools including Rust [6], Axum [2], Candle [4], and the broader open-source community for enabling this research.
+We thank the developers of open-source tools including Rust [klabnik2018rust], Axum [tokio2021axum], Candle [huggingface2023candle], and the broader open-source community for enabling this research.
+Declarations
+7.1 Funding
+Not applicable
+7.2 Conflict of interest/Competing interests
+Not applicable